Falconmaster:

If the sum of the interior angles of the hexagon at A, B, C, E and F is 595°, find the measure of the interior angle at D. A. 150° B. 210° C. 125° D. 60°

Falconmaster:

1 attachment
Angle:

this is when we can use the equation 180*(n - 2) = sum of interior angles

Falconmaster:

n is number of sides right?

Angle:

correct

Falconmaster:

so 6?

Angle:

yup

Falconmaster:

720

Angle:

Perfect then (sum of interior angles) = A + B + C + D + E + F but we are given that (A + B + C + E + F) = 595 so (sum of interior angles) = D + 595 solve for D

Falconmaster:

subtract the 500 or what not now?

Angle:

mhm

Falconmaster:

125?

Angle:

correct there you go

Falconmaster:

is that the answer i dont want to get it wrong XD

Angle:

yes it is this time

Falconmaster:

yay